OpenStudy (anonymous):

Find the area of the region bounded by the curves y2 = x, y – 4 = x, y = –2, and y = 1.

OpenStudy (anonymous):

help anyone? ive done these, but this is particularly hard.

OpenStudy (mathmale):

have you tried sketching the region described here?

OpenStudy (anonymous):

yes. the graph made it more confusing because there are 4 function but there are only 3 intersections. which i dont get because it doesnt create a "bounded" region.

OpenStudy (mathmale):

First I'll graph y=-2 and then y=1:

OpenStudy (mathmale):

Next: Do you recognize x=y-4? What does this represent (in terms of shape and orientation)?

OpenStudy (anonymous):

yes, that's a linear function

OpenStudy (mathmale):

Yes: x=y-4 is equivalent to y=x+4, as you could verify if you wanted to. Do you happen to have a graphing calculator? If so, are you familiar with its graphing capabilities?

OpenStudy (anonymous):